Actions
November 18-November 27th Plan¶
Scattering data with field, backup plan for straight tracks.
Daily schedule from 18/Nov to 27/Nov.¶
We expect to be running with 2V beamloss. Note Decay Solenoid is ON. ISIS is at 40+10 and 700 MeV
Date | Nov 18 | Nov. 19 | Nov. 20 | Nov. 21 | Nov. 22 |
MOM | E. Overton | E. Overton | E. Overton | E. Overton | E. Overton |
BLOC | Y.Song | Y.Song | Y.Song | J. Langlands | J. Langlands |
SOC | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram |
Lead shifter (08:00-16:00) | Y.Song | E. Overton | J. Martyniak | J. Ellison | J. Ellison |
Shifter 2 (08:00-16:00) | J.Ellison | J. Ellison | J. Ellison | H. Nebrensky | H. Nebrensky |
Lead shifter (16:00-00:00) | K. Long | F. Drielsma | Y. Song | Y. Song | F. Drielsma |
Shifter 2 (16:00-00:00) | A. Liu | A. Liu | F. Drielsma | F. Drielsma | A. Liu |
Magnet Monitor (00:00-08:00) | F. Drielsma | K. Long | A. Liu | A. Liu | Y. Song |
Shifter (00:00-08:00) | D. Ragaram | D. Ragaram + J Langlands | |||
Additional Personnel | P. Franchini | P. Franchini | |||
Notes | Morning training | ||||
Afternoon magnet ramp |
Date | Nov 23 | Nov. 24 | Nov. 25 | Nov. 26 | Nov. 27 |
MOM | E. Overton | E. Overton | E. Overton | E. Overton | E. Overton |
BLOC | J. Langlands | J. Langlands | J. Langlands | S. Boyd | S. Boyd |
SOC | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram | D. Ragaram |
Lead shifter (08:00-16:00) | V. Blackmore | J.B.Lagrange | J.B.Lagrange | S. Boyd | S. Boyd |
Shifter 2 (08:00-16:00) | J. Langlands | Y.Torn | Y. Torn | Y. Torn | Y. Torn |
Additional Personnel | |||||
Notes | |||||
Data to take¶
Goal: Take some useful data for the experiment. We have two possible scenarios, we have magnets, or we don't. The request list for each scenario is below..
Field On Data¶
Take task H1 first to give diagnostics data to check. Then move on to the official cooling data, start with task H2a, while settings for H2b,c are confirmed.
Task H1 (Done):¶
Test and optimise diffuser settings for 140MeV running
Task | Momentum | # Triggers | Beamline Tag | ColingChannel Tag | Done |
H1a | 140 | 50k at TOF2 | 3-140+M3-Test2 | 2016-04-01.3 | Y |
H1b | 140 | 50k at TOF2 | 6-140+M3-Test1 | 2016-04-01.3 | Y |
H1c | 140 | 50k at TOF2 | 10-140+M3-Test1 | 2016-04-01.3 | Y |
H1d | 140 | 10k at TOF2 | Ao Special | 2016-04-01.3 | 8577 |
Task H2 (Done):¶
Record data with 'official' settings
Task | Momentum | # Triggers | Beamline Tag | ColingChannel Tag | Similar runs |
H2a | 140 | 1,000k at TOF2 | 3-140+M3-Test2 | 2016-04-01.3 | ~400K TOF2 |
H2b | 140 | 500k at TOF1 | 10-140+M3-Test2 | 2016-04-01.3 | ~450K TOF1 |
H2c | 140 | 500k+ at TOF1 | 6-140+M3-Test2 | 2016-04-01.3 |
Task H3 (Done):¶
Muon beam for comparing with 7469
Task | Momentum | # Triggers | Beamline Tag | ColingChannel Tag | Similar runs |
H3 | 140 | 2 hours | 3-200+M0 | 2016-04-01.3 |
Tasks H4 + H5 moved to more recent run plan.¶
Straight track data (fallback)¶
Task S1: Detector alignment¶
Task | Momentum | # Triggers | Beamline Tag | ColingChannel Tag | Similar runs | Done |
S1a | 200 | 200k at TOF1 | 6-200+M0 | 8037 | Y | |
S1b | 300 | 1000k at TOF1 | 300MeV+pi pa82 | 8030 | Y | |
S1c | 400 | 200k at TOF1 | 400MeV+pi pa82 | 8041 | Y | |
S1d | 200 | 200k at TOF1 | 3-200+M3-Test1 | 8035 | Y |
Task S2: LiH scattering data - Moved to more recent plan.¶
TOF Calibration ? - Not needed/ Not done.¶
Todo:
+ Get settings from Durga
h2: Old daily schedules:
Updated by Overton, Edward almost 7 years ago ยท 51 revisions